The perpendicular OY is drawn from the centre O of the ellipse x2/a2 + y2/b2 = 1 to any tangent. Prove that the locus of Y is (x2 + y2 )2 = a2x2 + b2 y2.


1
Expert's answer
2021-08-19T07:02:36-0400

Let p(x,y) be the foot of perpendicular to a tangent

"y=mx\pm \sqrt{a^2m^2+b^2}\cdots (1)\\" from Centre

slope of OY = "\frac{-1}{m}\\"

Equation of OY is

"y=\frac{-1x}{m} \implies m=\frac{-x}{y} \cdots (2)"

p(x,y) lies on tangent,

"y=\frac{-x}{y}.x\pm\sqrt{a^2(\frac{-x}{y})^2+b^2}\\
y=\frac{-x}{y}.x\pm\sqrt{a^2\frac{x^2}{y^2}+b^2}\\
y=\frac{-x}{y}\pm\sqrt{\frac{a^2x^2+b^2y^2}{y^2}}\\
y^2+x^2=\pm\sqrt{a^2x^2+b^2y^2}\\
(x^2+y^2)^2=x^2a^2+b^2y^2."